# CERN-TH-2026-030

### v1

@@ Abstract
In this note, we give a non-perturbative construction of a lightlike domain wall separating IIA and IIB string theories in 10D in the framework of discrete light-cone quantization (DLCQ). In this setting, generalizations of the BFSS conjecture relate the 10D flat space limit to matrix string theories (MSTs) for IIA and IIB. The former is equivalent to the large-$N$ limit of 2D Super Yang-Mills theory, while the latter is the large-$N$ limit of 3D ABJM theory with $\pm 1$ Chern-Simons levels. Our construction requires the string coupling to vanish at the location of the wall, and we show that BPS IIA $D0$-branes become non-BPS IIB $D0$-branes as they cross the it, as anticipated in [1].
@@ 2 Brief Review of DLCQ of Type II String Theories
Not long after the BFSS model was introduced, it was realized that DLCQ constructions also follow for M-theory on $T^{n}$ at least when $n\leq 5$. In the limit of $\mathrm{Vol}(T^{n})\ll 1$ in 11D Planck units, this recovers IIA for $n=1$ and IIB for [^6] Recall that M-theory on a $T^{2}$ with vanishing area is equivalent to IIB where the $M2$ brane winding is identified with momentum along one of the spatial directions of 10D Minkowski space [25, 26, 27], a foundational fact in F-theory [28]. $n=2$. The analogs of the Super Quantum Mechanics of the NN $D0$ theory can be obtained by acting with T-duality. As we review below, the IIA and IIB 10D flat spacetime limits require an infinite gauge coupling limit of the 2D and 3D SYM theories respectively. The former recovers the usual Matrix String Theory [29, 18], while the latter can be shown to be equivalent to the CFT of $N\rightarrow\infty$ $M2$-branes, which is the ABJM theory at $k=1$. In what follows, we will simply call the former the IIA MST and the latter the IIB MST. [^7] To the author’s knowledge, this first proposal for the IIB MST appeared in [30, 31] well before ABJM. In fact one of the most recent references on this subject is [32] which appeared, incidentally, after the BLG-model [33, 34] on multiple $M2$-branes but still before ABJM. This means their method of treating a large-$N$ number of coincident $M2$-branes using BLG should, in retrospect, be replaced by $k=1$ ABJM CFTs at large-$N$ in light of [35, 36, 37].
@@ 3 IIA/IIB Wall
Figure 1: Matrix string theory construction of IIA/IIB wall. The $(-1)^{F_{L}}_{IIA}$ symmetry of the symmetric orbifold CFT in the IIA $g_{s}=0$ region is gauged to the right of the red line. The $g_{s}>0$ regions correspond to turning on certain irrelevant deformations of the 2D symmetric orbifold SCFTs. On the IIA side, this irrelevant deformation corresponds to moving flowing away from $g_{YM}^{2}=\infty$ to finite gauge coupling. On the IIB side, the irrelevant deformation corresponds to a 2D $\rightarrow$ 3D decompactification.
Definition: Let $g_{s}(\tau)=0$ for all $\tau$ the some neighborhood $(\tau_{0}-\epsilon,\tau_{0}+\epsilon)$, then one can form a lightlike IIA/IIB domain wall at $\tau_{0}$ by gauging the $(-1)^{F_{L}}_{IIA}$ symmetry (2.9) in the region $\tau\in[0,\tau_{0}+\epsilon)$. [^11] We will include the IIA and IIB subscripts on $(-1)^{F_{L}}$ unless otherwise obvious. [^12] Note that this is a symmetric orbifold analog of the half-space gauging considered on fundamental GS string worldsheets piercing the IIA/IIB wall considered in [1]. For early references on half-space gauging of discrete symmetries see [41, 42, 43]. See Figure 1 for an illustration of the MST setup.
This is well-defined in the limit $\epsilon\rightarrow 0$, where $g_{s}$ vanishes at an isolated value $\tau=\tau_{0}$, because along the free MST surface $\tau=\tau_{0}$, the half-interval gauging of $(-1)^{F_{L}}_{IIA}$ in $(\tau_{0}-\epsilon,\tau_{0}+\epsilon)$ reduces to a codimension-1 gauging at the $\tau=\tau_{0}$ line. Gauging a symmetry only along a codimension-1 submanifold is known as a codimension-$k$ condensation defect [43] which is a well-defined object for anomaly-free discrete symmetries. See Figure 2 for a spacetime illustration of this IIA/IIB wall located at the hypersurface $X^{+}=X^{+}_{0}$ ( $\tau=\tau_{0}$ in MST coordinates).
@@ 3.1 Objects Crossing the Wall
It was argued in [1] that BPS D-branes should become non-BPS D-branes when crossing the IIA/IIB wall. Here we study fate of IIA $D0$-branes crossing the IIA/IIB wall to understand if it indeed becomes a non-BPS $D0$-brane on the IIB side. In the language of MST, the $D0$-charge is $Q_{D0}=\int_{S^{1}_{R}}\mathrm{Tr}(F)$ [18]. This follows from the standard duality argument of IIA MST: the 2D $U(N)$ SYM describes NN $D1$s in a dual IIB frame and the flux $\mathrm{Tr}(F)$ sources D(1)D(-1)-branes which map to $D0$s under T-duality. In terms of the decomposition, $U(N)=(SU(N)\times U(1))/\mathbb{Z}_{N}$, a $Q_{D0}=1$ background corresponds to turning on a $1\;\mathrm{mod}\;N$ flux in the $SU(N)$ factor and a 1/N-1/N flux in the $U(1)$ factor. As explained in [18], such a flux sector has a diverging ground state energy in the $g_{s}\rightarrow 0$ limit, and indeed there is strong evidence that this $SU(N)$ flux vacuum is gapped [45]. To extract the open string spectrum of the $D0$-brane, we consider a modification of the original construction of the $D0$-brane in [18] to one localized along the $X^{-}$-direction. This can be done by sourcing the $Q_{D0}=1$ flux by a Wilson line at some $\sigma=\sigma_{A}$ in the $\mathbf{N}_{-1/N}$ representation of $(SU(N)\times U(1))/\mathbb{Z}_{N}$, as well as a Wilson line in the conjugate representation, $\overline{\mathbf{N}}_{+1/N}$, located at some $\sigma=\sigma_{B}$. See Figure 3 for an illustration. These Wilson lines are conformal interfaces in the $g_{s}\rightarrow 0$ limit and must be completely reflective in the sense of [46]. Therefore the choice of boundary conditions for the symmetric orbifold fields are either Dirichlet or Neumann. While both seem to in principle be allowed, we choose Dirichlet boundary conditions since we are interested in engineering the $D0$-brane massless modes. [^13] Other choices are most likely equivalent to considering bound states of various Type IIA BPS and non-BPS $D$-branes with the $D0$, but we leave this refined understanding of the IIA MST dictionary to future work. The $S_{N}$-invariant boundary conditions take the form [47]
$\partial_{+}X^{i}_{a}|_{\sigma=\sigma_{A,B}}=-\partial_{-}X^{i}_{a}|_{\sigma=\sigma_{A,B}},\quad\quad\theta^{\dot{\alpha}}_{a,+}|_{\sigma=\sigma_{A,B}}=\Gamma_{0}\theta^{\alpha}_{a,-}|_{\sigma=\sigma_{A,B}}$ (3.4)
@@
Figure 3: Symmetric orbifold configuration on $\mathbb{R}_{\tau}\times S^{1}_{\sigma}$ dual to a $D0$-brane localized along $X^{-}$. We have conformal interfaces (blue lines) separating the symmetric orbifold from the gapped $D0$-brane vacuum. The black lines are identified. The orange lines represent a single GS string with Dirichlet boundary conditions on the interfaces. These conformal interfaces are the IR limit of $U(N)$ Wilson lines in conjugate representations. Not pictured is are the overall gappless degrees of freedom in the $D0$ vacuum from the spin-0 and spin-1/2 components of the overall $U(1)$ vectormultiplet.
@@ 3.2 Spacetime Picture
$g_{s}(\tau)=\begin{cases}g_{s,0}\left(1-\exp(\frac{\tau^{2}}{\tau^{2}-c^{2}})\right),&|\tau|\leq c\\ 0,&|\tau|\geq c\end{cases}$ (3.8)
then the Ricci curvature very quickly grows to the order of string scale in the region $c\leq\tau\leq+c$. [^15] Note this parameter $c$ is not to be confused with $\epsilon$ from earlier. A IIA/IIB wall placed at $\tau\;(=X^{+})=0$ is can be described by Type II supergravity with constant dilaton/metric outside of this region, and requires the (irrelevant deformation of) the symmetric orbifold CFT description in the interior. This situation is not too dissimilar to the holographic dual solutions of $Dp$-branes worldvolume theories for $p\leq 2$ which always contain regions of the bulk spacetime which cannot be described by a gravitational EFT [23].
Finally, notice that because the half-space gauging interface is topological, the IIA/IIB wall is tensionless in the $g_{s}=0$ region. This means that its total mass-energy solely comes from the varying dilaton profile transverse to it which can only be calculated in the formalism of the symmetry orbifold picture. We leave such an analysis to future work.
@@ 4 Discussion and Outlook
In this note, we have given a non-perturbative definition of the IIA/IIB wall and initiated a study of its basic properties. It would be interesting to understand if similar constructions exist for $R7$-branes in Type II string theories. Indeed, it seems reasonable that a least light-like $R7$-branes can be constructed by defining the brane as a local operator in the MST theory which has a $(-1)^{F_{L}}$ topological symmetry line attached. Presumably, different choices of operators would correspond to different excited states of the $R7$. Matrix theory is a particularly appealing method to non-perturbatively define $R7$-branes since this is typically made difficult by the fact that well-studied AdS vacua often include fluxes which spontaneously break the gauge symmetry associated with the $R7$. For instance, the $F_{5}$-flux on Type IIB on $AdS_{5}\times S^{5}$ is not invariant under $(-1)^{F_{L}}$ nor $\Omega$, whereas even finite N MSTs have purely geometric holographic duals [21]. Similarly, given that there exists BFSS-like constructions for heterotic string theory [48, 49, 50, 51, 52, 53, 54, 55, 56], it would be interesting to pursue non-perturbative constructions of non-BPS branes of heterotic theories [6, 7].
@@
One interesting generalization is to consider non-BPS branes in asymmetric orbifold backgrounds which can also admit matrix theory constructions [58].

### v2

# arXiv:2603.02199v2[hep-th] 03 Mar 2026

@@ Abstract
In this note, we give a non-perturbative construction of a lightlike domain wall separating IIA and IIB string theories in 10D in the framework of discrete light-cone quantization (DLCQ). In this setting, generalizations of the BFSS conjecture relate the 10D flat space limit to matrix string theories (MSTs) for IIA and IIB. The former is equivalent to the large-$N$ limit of 2D Super Yang-Mills theory, while the latter is the large-$N$ limit of 3D ABJM theory with $\pm 1$ Chern-Simons levels. Our construction requires the string coupling to vanish at the location of the wall, and we show that BPS IIA $D0$-branes become non-BPS IIB $D0$-branes as they cross it, as anticipated in [1].
@@ 2 Brief Review of DLCQ of Type II String Theories
Not long after the BFSS model was introduced, it was realized that DLCQ constructions also follow for M-theory on $T^{n}$ at least when $n\leq 5$. In the limit of $\mathrm{Vol}(T^{n})\ll 1$ in 11D Planck units, this recovers IIA for $n=1$ and IIB for [^6] Recall that M-theory on a $T^{2}$ with vanishing area is equivalent to IIB where the $M2$ brane winding is identified with momentum along one of the spatial directions of 10D Minkowski space [25, 26, 27], a foundational fact in F-theory [28]. $n=2$. The analogs of the Super Quantum Mechanics of the NN $D0$ theory can be obtained by acting with T-duality. As we review below, the IIA and IIB 10D flat spacetime limits require an infinite gauge coupling limit of the 2D and 3D SYM theories respectively. The former recovers the usual Matrix String Theory [29, 18], while the latter can be shown to be equivalent to the CFT of $N\rightarrow\infty$ $M2$-branes, which is the ABJM theory at $k=1$. In what follows, we will simply call the former the IIA MST and the latter the IIB MST. [^7] To the author’s knowledge, the first proposal for the IIB MST appeared in [30, 31] well before ABJM. In fact one of the most recent references on this subject is [32] which appeared, incidentally, after the BLG-model [33, 34] on multiple $M2$-branes but still before ABJM. This means their method of treating a large-$N$ number of coincident $M2$-branes using BLG should, in retrospect, be replaced by $k=1$ ABJM CFTs at large-$N$ in light of [35, 36, 37].
@@ 3 IIA/IIB Wall
Figure 1: Matrix string theory construction of IIA/IIB wall. The $(-1)^{F_{L}}_{IIA}$ symmetry of the symmetric orbifold CFT in the IIA $g_{s}=0$ region is gauged to the right of the red line. The $g_{s}>0$ regions correspond to turning on certain irrelevant deformations of the 2D symmetric orbifold SCFTs. On the IIA side, this irrelevant deformation corresponds to flowing away from $g_{YM}^{2}=\infty$ to finite gauge coupling. On the IIB side, the irrelevant deformation corresponds to a 2D $\rightarrow$ 3D decompactification.
IIA/IIB Wall Definition: Let $g_{s}(\tau)=0$ for all $\tau$ the some neighborhood $(\tau_{0}-\epsilon,\tau_{0}+\epsilon)$, then one can form a lightlike IIA/IIB domain wall at $\tau_{0}$ by gauging the $(-1)^{F_{L}}_{IIA}$ symmetry (2.9) in the region $\tau\in[0,\tau_{0}+\epsilon)$. [^11] We will include the IIA and IIB subscripts on $(-1)^{F_{L}}$ unless otherwise obvious. [^12] Note that this is a symmetric orbifold analog of the half-space gauging considered on fundamental GS string worldsheets piercing the IIA/IIB wall considered in [1]. For early references on half-space gauging of discrete symmetries see [41, 42, 43]. See Figure 1 for an illustration of the MST setup. This is well-defined in the limit $\epsilon\rightarrow 0$, where $g_{s}$ vanishes at an isolated value $\tau=\tau_{0}$, because along the free MST surface $\tau=\tau_{0}$, the half-interval gauging of $(-1)^{F_{L}}_{IIA}$ in $(\tau_{0}-\epsilon,\tau_{0}+\epsilon)$ reduces to a codimension-1 gauging at the $\tau=\tau_{0}$ line. Gauging a symmetry only along a codimension-1 submanifold is known as a codimension-$k$ condensation defect [43] which is a well-defined object for anomaly-free discrete symmetries. See Figure 2 for a spacetime illustration of this IIA/IIB wall located at the hypersurface $X^{+}=X^{+}_{0}$ ($\tau=\tau_{0}$ in MST coordinates).
@@ 3.1 Objects Crossing the Wall
It was argued in [1] that BPS D-branes should become non-BPS D-branes when crossing the IIA/IIB wall. Here we study fate of IIA $D0$-branes crossing the IIA/IIB wall to understand if it indeed becomes a non-BPS $D0$-brane on the IIB side. In the language of MST, the $D0$-charge is $Q_{D0}=\int_{S^{1}_{R}}\mathrm{Tr}(F)$ [18]. This follows from the standard duality argument of IIA MST: the 2D $U(N)$ SYM describes NN $D1$s in a dual IIB frame and the flux $\mathrm{Tr}(F)$ sources D(1)D(-1)-branes which map to $D0$s under T-duality. In terms of the decomposition, $U(N)=(SU(N)\times U(1))/\mathbb{Z}_{N}$, a $Q_{D0}=1$ background corresponds to turning on a $1\;\mathrm{mod}\;N$ flux in the $SU(N)$ factor and a 1/N-1/N flux in the $U(1)$ factor. As explained in [18], such a flux sector has a diverging ground state energy in the $g_{s}\rightarrow 0$ limit, and indeed there is strong evidence that this $SU(N)$ flux vacuum is gapped [45]. To extract the open string spectrum of the $D0$-brane, we consider a modification of the original construction of the $D0$-brane in [18] to one localized along the $X^{-}$-direction. This can be done by sourcing the $Q_{D0}=1$ flux by a Wilson line at some $\sigma=\sigma^{\prime}$ in the $\mathbf{N}_{-1/N}$ representation of $(SU(N)\times U(1))/\mathbb{Z}_{N}$, as well as a Wilson line in the conjugate representation, $\overline{\mathbf{N}}_{+1/N}$, located at some $\sigma=\sigma^{\prime\prime}$. See Figure 3 for an illustration. These Wilson lines are conformal interfaces in the $g_{s}\rightarrow 0$ limit and must be completely reflective in the sense of [46]. Therefore the choice of boundary conditions for the symmetric orbifold fields are either Dirichlet or Neumann. While both seem to in principle be allowed, we choose Dirichlet boundary conditions since we are interested in engineering the $D0$-brane massless modes. [^13] Other choices are most likely equivalent to considering bound states of various Type IIA BPS and non-BPS $D$-branes with the $D0$, but we leave this refined understanding of the IIA MST dictionary to future work. The $S_{N}$-invariant boundary conditions take the form [47]
$\partial_{+}X^{i}_{a}|_{\sigma=\sigma^{\prime},\sigma^{\prime\prime}}=-\partial_{-}X^{i}_{a}|_{\sigma=\sigma^{\prime},\sigma^{\prime\prime}},\quad\quad\theta^{\dot{\alpha}}_{a,+}|_{\sigma=\sigma^{\prime},\sigma^{\prime\prime}}=\Gamma_{0}\theta^{\alpha}_{a,-}|_{\sigma=\sigma^{\prime},\sigma^{\prime\prime}}$ (3.4)
@@
Figure 3: Symmetric orbifold configuration on $\mathbb{R}_{\tau}\times S^{1}_{\sigma}$ dual to a BPS IIA $D0$-brane localized along $X^{-}$. We have conformal interfaces (blue lines) separating the symmetric orbifold from the gapped $D0$-brane vacuum. The black lines are identified. The orange lines represent a single GS string with Dirichlet boundary conditions on the interfaces. These conformal interfaces are the IR limit of $U(N)$ Wilson lines in conjugate representations. Not pictured is are the overall gappless degrees of freedom in the $D0$ vacuum from the spin-0 and spin-1/2 components of the overall $U(1)$ vectormultiplet.
@@ 3.2 Spacetime Picture
$g_{s}(\tau)=\begin{cases}g_{s,0}\left(1-\exp(\frac{\tau^{2}}{\tau^{2}-c^{2}})\right)&|\tau|\leq c\\ g_{s,0}&|\tau|\geq c\end{cases}$ (3.8)
then the Ricci curvature very quickly grows to the order of string scale in the region $c\leq\tau\leq+c$. [^15] Note that this parameter $c$ is not to be confused with $\epsilon$ from earlier. A IIA/IIB wall placed at $\tau\;(=X^{+})=0$ is can be described by Type II supergravity with constant dilaton/metric outside of this region, and requires the (irrelevant deformation of) the symmetric orbifold CFT description in the interior. This situation is not too dissimilar to the holographic dual solutions of $Dp$-branes worldvolume theories for $p\leq 2$ which always contain regions of the bulk spacetime which cannot be described by a gravitational EFT [23].
Finite string frame tension:
Finally, notice that because the half-space gauging interface is topological, the IIA/IIB wall is tensionless in the $g_{s}=0$ core region. This means that its total mass-energy comes solely from the varying dilaton profile transverse to it which can only be calculated near the core region in the symmetric orbifold formalism due to the breakdown of the Type II supergravity EFT when $|\nabla\phi|$ is large. This allows us to conclude that the tension of our IIA/IIB wall must be finite in the string frame [^16] The string frame tension is directly related to the MST energy because the MST Hamiltonian is directly proportional to the string worldsheet Hamiltonian when $g_{s}$ is small [18]. because $g_{s}(\tau)=0$ is adiabatically connected to the case of a small position-dependent $g_{s}(\tau)$, such as $g_{s}(\tau)=g_{s,0}(1+\tau^{2}/c^{2})$ in the neighborhood of $\tau=0$ in (3.8). This follows from the symmetric orbifold point-of-view since this just amounts to turning on a small position dependent irrelevant coupling. This implies that the string frame tension of the IIA/IIB wall is adiabatically connected to zero, so therefore must be finite. While we leave a precise calculation of this tension as a function(al) of $g_{s}(\tau)$ for future work, we can conclude that our domain wall satisfies the Swampland Cobordism Conjecture prediction [8] of a finite energy domain wall between IIA and IIB since the $\tau$-dependent MST Hamiltonian takes $\mathrm{Sym}^{N}(\mathbb{R}^{8})$ states to $\mathrm{Sym}^{N}(\mathbb{R}^{8})/\mathbb{Z}_{2}$ states (with small irrelevant couplings on both sides) in finite time. Additionally, for each value of $N$, such a 2D/3D gauge theory interface precisely engineers a cobordism class-killing domain wall in an AdS holographic setting [48].
@@ 4 Discussion and Outlook
In this note, we have given a non-perturbative definition of the IIA/IIB wall and initiated a study of its basic properties. It would be interesting to understand if similar constructions exist for $R7$-branes in Type II string theories. Indeed, it seems reasonable that a least light-like $R7$-branes can be constructed by defining the brane as a local operator in the MST theory which has a $(-1)^{F_{L}}$ topological symmetry line attached. Presumably, different choices of operators would correspond to different excited states of the $R7$. Matrix theory is a particularly appealing method to non-perturbatively define $R7$-branes since this is typically made difficult by the fact that well-studied AdS vacua often include fluxes which spontaneously break the gauge symmetry associated with the $R7$. For instance, the $F_{5}$-flux on Type IIB on $AdS_{5}\times S^{5}$ is not invariant under $(-1)^{F_{L}}$ nor $\Omega$, whereas even finite $N$ MSTs have purely geometric holographic duals [21]. It would be interesting to make contact with the $R7$-brane gravitational solutions found recently in [49].
Given that there exists BFSS-like constructions for heterotic string theory [50, 51, 52, 53, 54, 55, 56, 57, 58], it would also be interesting to pursue non-perturbative constructions of non-BPS branes of heterotic theories [6, 7]. Also, one may consider non-BPS branes in asymmetric orbifold backgrounds as these also admit matrix theory constructions [59].